# Enhancing Arabic Automated Essay Scoring with Synthetic Data and Error Injection

### v1

@@ 1 Introduction
Automated Essay Scoring (AES) is a technology that automates the evaluation and scoring of essays to assess language learners’ writing quality while eliminating the need for human intervention Shermis and Burstein (2003). AES has gained great interest due to its significant benefits in the field of education Lagakis and Demetriadis (2021); Susanti et al. (2023). AES systems help teachers evaluate many essays with consistent scoring and reduced workload. On the other hand, AES helps students improve their writing quality through rapid real-time scoring and feedback Hahn et al. (2021). Unlike for English, for Modern Standard Arabic (MSA) it is difficult to develop robust and scalable AES systems, primarily due to the lack of essay datasets necessary for building effective Arabic AES Lim et al. (2021); Elhaddadi et al. (2024). This paper presents our proposed framework to tackle the issue of data scarcity and quality by utilizing Transformers and Large Language Models (LLMs) to generate and build a synthetic dataset. Our approach begins with prompting GPT-4o to generate a variety of Arabic essays covering multiple topics and different writing proficiency levels as defined by the Common European Framework of Reference (CEFR) Council of Europe (2001). Subsequently, we use a controlled error injection model to introduce errors into the correct Arabic essays, ensuring that erroneous essays reflect the type of errors that are commonly made by learners of Arabic in real-world scenarios. Our error injection approach consists of two steps: (i) Error Type Prediction, where a fine-tuned CAMeLBERT MSA model Inoue et al. (2021) classifies the most likely error type for each word, and (ii) Error Realization, where we apply a bigram MLE model to determine the most probable transformation for each predicted error type. Our framework enables the generation of realistic human-like essays, enhancing data augmentation for Arabic AES systems. Our main contributions are as follows:
@@
  • Creating a synthetic Arabic AES dataset with 3,040 essays annotated with CEFR proficiency levels. [^1] The full dataset will be publicly available for research.
  • Developing an Arabic essay auto-marking system using a BERT-based scoring model, enabling accurate and scalable evaluation of Arabic essays based on CEFR standards.
The rest of the paper is organized as follows: §2 reviews related work on AES, §3 describes the dataset, and §4 outlines our data augmentation approach. §5 details the error injection methods, followed by evaluation in §6 and §7 presents conclusion and future work.
@@ 2 Related Work
The field of AES has been investigated extensively, particularly in English Lim et al. (2021); Ramesh and Sanampudi (2022), where some valuable tools have been introduced such as IntelliMetric Elliott et al. (2003), e-rater Attali and Burstein (2006), Grammarly, [^2] https://app.grammarly.com/ Write and Improve [^3] https://writeandimprove.com/ Yannakoudakis et al. (2018), and others. The development and validation of English AES have been enabled by the large-scale annotated datasets that have been released, such as the First Cambridge English (FCE) dataset [^4] https://ilexir.co.uk/datasets/index.html Yannakoudakis et al. (2011), Automated Student Assessment Prize (ASAP) dataset, [^5] https://www.kaggle.com/c/asap-aes the TOEFL11 corpus Blanchard (2013), and the ICLE (International Corpus of Learner English) Granger (2003). These datasets contained thousands of student essays with proficiency-level grades, often along multiple dimensions. In contrast, research on Arabic AES has not received comparable attention. Some studies have focused on Arabic AES, utilizing feature engineering and applying machine learning to develop models Alghamdi et al. (2014); Al-Shalabi (2016); Alobed et al. (2021); Gaheen et al. (2021). This work and related research has only partially addressed the key challenges hindering the development of Arabic AES models, namely, the lack of publicly available large-scale annotated datasets designed to improve the quality of Arabic writing. Ghazawi and Simpson (2024) introduced AR-AES, a benchmark dataset that contains 2,046 essays written by undergraduate students across three different university faculties, featuring a mix of question types. Two educators from each faculty annotated and graded the essays using clearly defined rubrics, aiming to assess students’ academic performance across various disciplines. Unlike AR-AES, our work aims to evaluate and assess writing proficiency using the CEFR rubrics as these constitute an internationally accepted standard. QAES Bashendy et al. (2024) presented the first publicly available trait-specific annotations for Arabic AES. QAES is based on the work of the Qatari Corpus of Argumentative Writing (QCAW) Ahmed et al. (2024), which consists of 195 Arabic argumentative essays. They proposed a multi-layer of annotation traints such as coherence, organization, grammar, and others. Despite its excellent annotation, it is small in size and limited to two prompts. While QAES multi-traits scores are publicly available, the QCAW holistic score is not public. Habash and Palfreyman (2022) presented the Zayed University Arabic-English Bilingual Undergraduate Corpus (ZAEBUC). This corpus comprises non-parallel essays in Arabic and English related to three prompts collected from first-year university students with differing writing proficiency. ZAEBUC includes 216 annotated Arabic essays featuring manual annotations for syntactic and morphological characteristics and a CEFR-based proficiency assessment. Again, ZAEBUC is small in size and limited to three prompts. Researchers have investigated several approaches to data augmentation such as sampling techniques, introducing noise into texts, or paraphrasing, to overcome the challenges of data scarcity and quality Li et al. (2022). The recent development of LLMs has paved the way for researchers to explore promising new data synthesis solutions Wang et al. (2024a); Long et al. (2024). Transformers and LLMs can closely mirror real-world distributions while introducing valuable variations across multiple tasks and domains Wang et al. (2024b). GPT models such as ChatGPT, GPT-4, and GPT-4o demonstrate significant capabilities in augmenting and generating synthetic essays for the purpose of English AES Ramesh and Sanampudi (2022). LLMs and Transformer models with Arabic capabilities have shown robust generalization across different Arabic NLP tasks such as: Question-Answering Samuel et al. (2024), Code-Switching Alharbi et al. (2024), Name Entity Recognition (NER) Sabty et al. (2021), Grammar Error Correction (GEC) Solyman et al. (2023), and Sentiment Analysis Refai et al. (2023). However, to the best of our knowledge no research has utilized such models to generate Arabic essays across CEFR writing proficiency levels.
@@
3 Data: The ZAEBUC Corpus
For all our experiments and reported results, we use the ZAEBUC corpus Habash and Palfreyman (2022). ZAEBUC comprises essays written by native Arabic speakers, which were manually corrected and annotated for writing proficiency using the CEFR Council of Europe (2001) rubrics and scale. Each essay was independently annotated by three CEFR-proficient bilingual speakers. Habash and Palfreyman (2022), assigned a holistic CEFR level to each essay by converting the three CEFR ratings into numerical scores (ranging from 1 to 6) and then taking the rounded average. The essays in the corpus were limited to three prompt choices on Social Media, Tolerance, and Development; see Table 1. We use the splits created by Alhafni et al. (2023). Table 2 presents the ZAEBUC corpus CEFR level distributions in terms of the holistic CEFR levels.
@@ 4.1 Building Essay Prompts
We began by compiling a collection of essay prompts covering diverse categories and proficiency levels. We considered three proficiency levels aligned with the CEFR scale: Beginner (A1–A2), Intermediate (B1–B2), and Advanced (C1–C2). While some categories, such as Hobbies, are suitable for all levels, others, like Politics, Technology, and Education, were introduced specifically for advanced learners. Using various LLMs, including GPT-4o, Gemini, and Copilot, we generated 100 prompts across these categories, ensuring relevance for Arabic essay writing and an unbiased distribution across proficiency levels. We then manually reviewed the prompts, removing redundancies and irrelevant entries while maintaining comprehensive coverage. The final collection consists of 152 balanced and diverse prompts. Table 3 presents the selected categories and the distribution of prompts across levels, while Table 4 provides example prompts from the Hobbies category.
@@ 4.3 Zero-shot Data Augmentation
Effective LLM prompt engineering has become increasingly important, as the model’s output varies based on the prompt, provided instructions, and prompt language. Previous studies in Arabic NLP have shown that using English as the instruction language for input prompts can improve output quality Kmainasi et al. (2024); Koto et al. (2024). In our approach, we experiment with various prompts for zero-shot data augmentation to identify those that produce human-like text while adhering to guideline instructions. We use GPT-4o as our generation model due to its affordability and larger token capacity for both input and output. The GPT prompts include: (a) the target CEFR level, (b) CEFR guidelines and instructions, (c) the linguistic profile for the targeted CEFR level to control the prompt output, and (d) the topic prompt or question from the previously mentioned topic prompts dataset. To check the quality of the generated essays and whether they follow the prompt instructions, we build a linguistic feature profile for each augmented essay. We employ the cosine similarity to assess the faithfulness of the augmented approach and whether the augmented essays are close to the original CEFR profiles. We applied the cosine similarity as in equation 3 either to one of the features profiles (lexcial/syntactical features , morphological features or both of them).
@@
where $P_{i}$ and $Q_{i}$ are the probability values of linguistic features in the original and generated distribution profiles. This metric ensures that the synthetic data closely aligns with real human essay patterns. Later we calculate the matching agreements by:
@@
where $\hat{y}_{i}$ is the predicted level and $y_{i}$ is the ground truth. We conducted multiple rounds of prompt engineering refinements to improve the quality of the generated Arabic essays and ensure alignment with CEFR levels. First, we found that straightforward prompts without explicit controlled linguistic instructions and explanations resulted in incoherent essays, including irrelevant topics and English text, achieving only 20.5% matching agreement with linguistic feature profiles. In a subsequent round, we introduced detailed definitions of linguistic features and restricted outputs to Arabic-only text, which improved agreement to 26%. However, the model still occasionally produced incomplete essays and injected text from the prompt into the essay. The most effective prompt structure format is illustrated in Figure 1. We separated system-level control instructions from user-defined parameters, thereby providing clearer guidance for structured and proficiency-aligned text generation. This refinement increased agreement to 27.5%, demonstrating that precisely controlled instructions enhance LLM performance in structured writing tasks. Ultimately, we generated 3,040 Arabic essays covering all CEFR levels and various topics, where each prompt produced ten essays. The structured and controlled prompt design not only improved alignment with learner writing styles but also established a consistent framework for synthetic Arabic text generation. Table 5 presents the distribution of generated essays across different CEFR levels.
@@
Figure 1: GPT-4o prompts messages that has been used to generate Arabic essays
@@ 5 Error Injection
Human-generated text naturally contains some grammatical errors and linguistic infelicities. In order to generate human-like essays, we need to include similar kinds of errors that reflect the level of writing attainment. In this phase, we prompt GPT-4o to inject errors into the previously generated essays while maintaining their aligned CEFR levels by utilizing error profiling.
@@ 5.1 Error Profiling
To model the distribution of errors to inject into the synthetic essays, we again leverage the ZAEBUC corpus, which contains both the erroneous essays aligned with the manually corrected ones. We followed the same methodology we used to construct the linguistic feature profiles for each CEFR level to develop error distribution profiles aligned with CEFR levels. The error profile captures and reflects the authentic distribution patterns observed in human writing at different CEFR levels. Developing an Error Instruction Repository To prompt GPT-4o to generate essays containing errors we applied the GED model proposed by Alhafni et al. (2023) to the ZAEBUC corpus to annotate errors using 13 error tags and to obtain error distributions for each CEFR level. We created the repository using the same 13 tags (ReplaceM, ReplaceO, ..), where we also added a formal definition of what those tags describe in terms of linguistic errors. In addition, we expanded the error taxonomy by splitting it into finer-grained classes. Each error instruction is followed by an example showing the correct word and the erroneous version. The explanation was based on the extended ALC taxonomy Alfaifi et al. (2013), which was refined later and introduced as ARETA Belkebir and Habash (2021). Figure 2 shows some examples from our error instruction repository.
@@ 5.2 GPT-Based Error Injection
We prompted GPT-4o to inject errors into the synthetically generated essays based on the error distribution profiles while maintaining the CEFR level. The model processed one essay at a time in a zero-shot setting, except that we included the definition and explanation of the error tags. For example, Replace_M indicates a morphological error, while Merge targets two mistakenly split tokens that need to be merged, and so on. After conducting multiple experiments, we observed the following issues: (i) The model struggled to follow the predefined error distribution perhaps due to the complexity of the prompts. (ii) The model was confused by certain error tags, particularly SPLIT, MERGE as well as the UNK (unknown errors). These errors were mainly ignored in the injected text. (iii) We calculated the cosine similarity between the main error profile and the injected essays’ error distributions as shown in equation 3. When we injected all errors at once, the similarity agreements did not exceed 20%; however, when we reduced the number of error tags per essay or omitted the misleading tags, the agreements significantly improved, reaching 86% Therefore we implemented a method where each error type was injected separately. This required multiple iterations over the same essay, corresponding to the number of error tags shown in the error distribution profile for each CEFR level. Figure 3 shows an example of a GPT-4o prompt for error injection. Some errors occurred more frequently than others among Arabic writers. For example, orthographic errors are more frequent than semantic errors. So, rather than repeating a single error instruction for all instances of an error tag, we randomly select a weighted number of error instructions proportional to the average occurrence of each error type. Figure 4 shows the pseudocode for the selection process. The full pseudocode is in Appendix A.

### v2

@@ 1 Introduction
Automated Essay Scoring (AES) is a technology that automates the evaluation and scoring of essays to assess language learners’ writing quality while eliminating the need for human intervention Shermis and Burstein (2003). AES has gained great interest due to its significant benefits in the field of education Lagakis and Demetriadis (2021); Susanti et al. (2023). AES systems help teachers evaluate many essays with consistent scoring and reduced workload. On the other hand, AES helps students improve their writing quality through rapid real-time scoring and feedback Hahn et al. (2021). Unlike for English, it is difficult to develop robust and scalable AES systems for Modern Standard Arabic (MSA), primarily due to the lack of essay datasets necessary for building effective Arabic AES Lim et al. (2021); Elhaddadi et al. (2024). This paper presents a framework to tackle the issue of data scarcity and quality by utilizing Transformers and Large Language Models (LLMs) to generate and build a synthetic dataset. Our approach begins with prompting GPT-4o to generate a variety of Arabic essays covering multiple topics and different writing proficiency levels as defined by the Common European Framework of Reference (CEFR) Council of Europe (2001). Subsequently, we use a controlled error injection model to introduce errors into the correct Arabic essays, ensuring that erroneous essays reflect the type of errors that are commonly made by learners of Arabic in real-world scenarios. Our error injection approach consists of two steps: (i) Error Type Prediction, where a fine-tuned CAMeLBERT MSA model Inoue et al. (2021) classifies the most likely error type for each word, and (ii) Error Realization, where we apply a bigram MLE model to determine the most probable transformation for each predicted error type. Our framework enables the generation of realistic human-like essays, enhancing data augmentation for Arabic AES systems. Our main contributions are as follows:
@@
  • Creating a synthetic Arabic AES dataset with 3,040 essays annotated with CEFR proficiency levels.
  • Developing an Arabic AES system using a BERT-based model, enabling accurate and scalable evaluation of Arabic essays based on CEFR standards.
The rest of the paper is organised as follows: §2 reviews related work on AES, §3 describes the dataset, and §4 outlines our data augmentation approach. §5 details the error injection methods, followed by an evaluation in §6. We discuss our results in §7 and §8 presents the conclusion and future work.
@@ 2 Related Work
AES has been investigated extensively, particularly in English Lim et al. (2021); Ramesh and Sanampudi (2022), where multiple tools have been introduced such as IntelliMetric Elliott et al. (2003), e-rater Attali and Burstein (2006), Grammarly, [^2] https://app.grammarly.com/ Write and Improve [^3] https://writeandimprove.com/ Yannakoudakis et al. (2018), and others. The development of English AES systems has been enabled by large scale annotated datasets such as the First Cambridge English (FCE) dataset [^4] https://ilexir.co.uk/datasets/index.html Yannakoudakis et al. (2011), Automated Student Assessment Prize (ASAP) dataset, [^5] https://www.kaggle.com/c/asap-aes the TOEFL11 corpus Blanchard (2013), and the ICLE (International Corpus of Learner English) Granger (2003). These datasets contain thousands of student essays with proficiency level grades, often along multiple dimensions. In contrast, Arabic AES research has received less attention. Some studies have applied feature engineering and machine learning to develop models Alghamdi et al. (2014); Al-Shalabi (2016); Alobed et al. (2021); Gaheen et al. (2021), but they partially address key challenges, especially the scarcity of large, publicly available annotated datasets for improving Arabic writing quality. Ghazawi and Simpson (2024) introduced AR-AES, a benchmark of 2,046 undergraduate essays from three university faculties, annotated by two educators per faculty using rubrics to assess academic performance. In contrast, our work focuses on writing proficiency, using the CEFR standard. Bashendy et al. (2024) presented QAES, the first publicly available trait specific annotations for Arabic AES. QAES extends the Qatari Corpus of Argumentative Writing (QCAW) Ahmed et al. (2024), which consists of 195 Arabic argumentative essays. They implemented multi-layered annotation of traits such as coherence, organization, grammar, and others. Despite its comprehensive annotation, it is small in size and limited to two prompts. While QAES multi-traits scores are publicly available, the QCAW holistic score is not. Habash and Palfreyman (2022) presented the Zayed University Arabic-English Bilingual Undergraduate Corpus (ZAEBUC). This corpus comprises non-parallel essays in Arabic and English related to three prompts collected from first-year university students with differing writing proficiency. ZAEBUC includes 216 annotated Arabic essays featuring manual annotations for syntactic and morphological characteristics and a CEFR-based proficiency assessment. Again, ZAEBUC is small in size and limited to three prompts. Researchers have explored data augmentation methods like sampling, noise injection, and paraphrasing to address data scarcity and quality Li et al. (2022). The recent development of LLMs has paved the way for researchers to explore promising new data synthesis solutions Wang et al. (2024a); Long et al. (2024). Transformers and LLMs can closely mirror real-world distributions while introducing valuable variations across multiple tasks and domains Wang et al. (2024b). GPT models have shown strong capabilities in generating synthetic essays for English AES Ramesh and Sanampudi (2022). LLMs and Transformers have also generalized well in Arabic NLP tasks, including Question Answering Samuel et al. (2024), Code Switching Alharbi et al. (2024), NER Sabty et al. (2021), Grammatical Error Correction Alhafni and Habash (2025); Solyman et al. (2023), and Sentiment Analysis Refai et al. (2023). However, to the best of our knowledge no research has utilized such models to generate Arabic essays across CEFR writing proficiency levels.
3 Data: The ZAEBUC Corpus
For all our experiments, we use the ZAEBUC corpus Habash and Palfreyman (2022). ZAEBUC comprises essays written by native Arabic speakers, which were manually corrected and annotated for writing proficiency using the CEFR Council of Europe (2001) rubrics and scale. Each essay was annotated by three CEFR-proficient bilingual speakers. Habash and Palfreyman (2022), assigned a holistic CEFR level to each essay by converting the three CEFR ratings into numerical scores (ranging from 1 to 6) and then taking the rounded average. The essays in the corpus were limited to three prompt choices on Social Media, Tolerance, and Development; see Table 1. We use the splits created by Alhafni et al. (2023). Table 2 shows the CEFR level distribution of the ZAEBUC corpus based on holistic CEFR scores. The ZAEBUC corpus is limited in size and skewed toward B1–B2 levels, with no A1 or C2 essays. This common imbalance in Arabic learner data motivated our synthetic approach to create a more balanced CEFR distribution.
@@ 4.1 Building Essay Prompts
We began by compiling a diverse set of essay prompts across various categories and CEFR levels. While not directly drawn from established frameworks, our prompts were inspired by themes common in language assessments, including placement tests and academic writing. We aimed to cover familiar and level-appropriate topics, such as social issues, education, and personal experiences, while ensuring balance across the CEFR bands. We considered three proficiency levels: Beginner (A1–A2), Intermediate (B1–B2), and Advanced (C1–C2). General themes, such as hobbies, suited all levels, while more complex topics, including politics, Technology, and Education, were reserved for advanced learners. Using LLMs like GPT-4o [^6] https://openai.com/index/hello-gpt-4o/, Gemini [^7] https://gemini.google.com/app, and Copilot [^8] https://copilot.microsoft.com/, we generated 100 prompts, followed by a manual review to remove redundancies and ensure both relevance for Arabic essay writing and balanced proficiency coverage. The final collection consists of 152 balanced and diverse prompts. Table 3 presents the selected categories and the distribution of the prompts across levels, while Table 4 provides example prompts for the Hobbies category.
@@ 4.3 Zero-shot Data Augmentation
Effective LLM prompt engineering has become increasingly important, as the model’s output varies based on the prompt, provided instructions, and prompt language. Previous studies in Arabic NLP have shown that using English as the instruction language for input prompts can improve output quality Kmainasi et al. (2024); Koto et al. (2024). In our approach, we experiment with various prompts for zero-shot data augmentation to identify those that produce human-like text while adhering to guideline instructions. We use GPT-4o as our generation model due to its affordability and larger token capacity for both input and output. The GPT prompts include (a) the target CEFR level, (b) CEFR guidelines and instructions, (c) the linguistic profile for the targeted CEFR level to control the prompt output, and (d) the topic prompt or question from the previously mentioned topic prompts dataset. For these missing levels (A1 and C2), instead of injecting a pre-defined profile, GPT-4o was directly prompted to act as an assistant and generate data based on the general standards and rubrics of the CEFR. To check the quality of the generated essays and whether they follow the prompt instructions, we build a linguistic feature profile (vector) for each augmented essay. We then assess the alignment between the generated essays and the reference CEFR-level profiles by computing their feature vectors’ cosine similarity as in equation 3. Specifically, given two real-valued feature vectors $P_{i}$ (the CEFR reference profile) and $Q_{i}$(the generated essay), the cosine similarity is calculated as:
@@
This metric ensures that the synthetic data closely aligns with real human essay patterns. Based on the computed similarity score, we assign a predicted CEFR level to each essay. Later, we calculate the alignment between the predicted CEFR level and the target level specified in the GPT-4o prompt (ground truth) using the following agreement formula:
@@
where $\hat{y}_{i}$ is the predicted level and $y_{i}$ is the ground truth. This process evaluates how well GPT-4o succeeded in aligning the generated content with the intended proficiency level, serving as a measure of agreement rather than a prediction from an external model. We conducted multiple rounds of prompt engineering refinements to improve the quality of the generated Arabic essays and ensure alignment with CEFR levels. First, we found that straightforward prompts without explicit controlled linguistic instructions and explanations resulted in incoherent essays, including irrelevant topics and English text, achieving only 20.5% matching agreement with linguistic feature profiles. In a subsequent round, we introduced detailed definitions of linguistic features and restricted outputs to Arabic-only text, which improved agreement to 26%. However, the model still occasionally produced incomplete essays and injected text from the prompt into the essay. The most effective prompt structure format is illustrated in Figure 1. We separated system-level control instructions from user-defined parameters, thereby providing clearer guidance for structured and proficiency-aligned text generation. This refinement increased agreement to 27.5%, demonstrating that precisely controlled instructions enhance LLM performance in structured writing tasks. Ultimately, we generated 3,040 Arabic essays covering all CEFR levels and various topics, where each prompt was used to create ten essays. This effort was intentionally designed to address the imbalanced CEFR distribution in the original ZAEBUC corpus, where B-level essays were overrepresented. By constructing a more balanced synthetic dataset, we aimed to enhance model performance across the full proficiency spectrum. The structured and controlled prompt design also improved alignment with learner writing styles while providing a consistent framework for generating realistic Arabic essays. Table 5 presents the distribution of generated essays across different CEFR levels. The full dataset statistics are provided in Appendix A.1.
@@
Figure 1: GPT-4o prompts messages that have been used to generate Arabic essays
@@ 5 Error Injection
Human-generated text naturally contains some grammatical errors and linguistic infelicities. In order to create human-like essays, we need to add similar kinds of errors to the synthetic essays that reflect the level of writing attainment. In this phase, we prompt GPT-4o to inject errors into the previously generated essays while maintaining their aligned CEFR levels by utilizing error profiling.
@@ 5.1 Error Profiling
To model the distribution of errors to inject into the synthetic essays, we again leverage the ZAEBUC corpus, which contains the erroneous essays aligned with the manually corrected ones. We followed the same methodology we used to construct the linguistic feature profiles for each CEFR level to develop error distribution profiles aligned with CEFR levels. The error profile captures and reflects the authentic distribution patterns observed in human writing at different CEFR levels. Developing an Error Instruction Repository To prompt GPT-4o to generate essays containing errors we applied the Grammatical Error Detection (GED) model proposed by Alhafni et al. (2023) to the ZAEBUC corpus to annotate errors using 13 error tags and to obtain error distributions for each CEFR level. We created the repository using the error tags, where we also added a formal definition of what those tags describe in terms of linguistic errors. In addition, we expanded the error taxonomy by splitting it into finer-grained classes. Each error instruction is followed by an example showing the correct word and the erroneous version. The explanation was based on the extended ALC taxonomy Alfaifi et al. (2013), which was refined later and introduced as ARETA Belkebir and Habash (2021). Appendix C presents examples of the error types. Figure 2 shows some examples from our error instruction repository.
@@ 5.2 GPT-Based Error Injection
We prompted GPT-4o to inject errors into the synthetically generated essays based on the error distribution profiles while maintaining the CEFR level. The model processed one essay at a time in a zero-shot setting, except that we included the definition and explanation of the error tags. For example, M indicates a morphological error, while Merge targets two mistakenly split tokens that need to be merged, and so on. After conducting multiple experiments, we observed the following issues: (i) The model struggled to follow the predefined error distribution perhaps due to the complexity of the prompts. (ii) The model was confused by certain error tags, particularly Split and Merge. These errors were mainly ignored in the injected text. (iii) We calculated the cosine similarity between the main error profile and the injected essays’ error distributions as shown in Equation 3. When we injected all errors at once, the similarity agreements did not exceed 20%; however, when we reduced the number of error tags per essay the agreements significantly improved, reaching 86%. Therefore we implemented a method where each error type was injected separately. This required multiple iterations over the same essay, corresponding to the number of error tags shown in the error distribution profile for each CEFR level. Figure 3 shows an example of a GPT-4o prompt for error injection. Some error types, especially orthographic errors, are more frequent among Arabic writers than others. The prompt was intentionally designed through prompt engineering. The ‘helpful assistant’ component establishes a cooperative persona for the LLM, while the subsequent instruction to ‘inject erroneous tokens’ explicitly guide GPT-4o towards the specific task of error introduction. This approach ensures that GPT-4o is not making random edits but is rather following predefined instructions to create targeted errors, aligning with the overall goal of generating realistic synthetic data. To reflect this, we randomly select weighted error instructions based on the average frequency of each error type. Figure 4 shows the pseudocode for the selection process. The full pseudocode is in Appendix B.